OpenStudy (anonymous):

Complete the table and choose the solution for each given value of x in questions 8–13. y = –x2 + 2

OpenStudy (anonymous):

When x = –2

OpenStudy (anonymous):

HELP!

OpenStudy (kropot72):

\[y=-x ^{2}+2\] When x = -2\[x ^{2}=-2\times -2=4\] y = -4 + 2 = -2

OpenStudy (anonymous):

Thaannkkk yoouu!! :)

OpenStudy (kropot72):

You're welcome :)